OpenStudy (anonymous):

translate the problem situation to a system of equations. do not solve. bill & jose's discount cinema sold adult tickets for $2.80 each and children's tickets for $1.80 each. last Tuesday, a total of $71.80 was collected from 36 movie watchers. how many of each type were sold

OpenStudy (anonymous):

x + y = 4.60; 2.80x + 1.80y = $71.80 In this case you've assigned children as x, and adults as y. You must write the second equation as: x+y=36 because there were 36 movie watchers (adults and children). Now if you solve this system, you get the following: 7 children and 29 adults.